OpenStudy (anonymous):

Find the ratio of the slant height to the radius of a right circular cone whose volume and lateral area are numerically equal. Assume the altitude of the cone to be 9 units

OpenStudy (anonymous):

let r be radius l be slant height,h be height of cone. \[volume V=\frac{ 1 }{ 3 }\pi r^2 h\] \[lateral ~ area~A=\pi rl\] h=9 \[V=\frac{ 1 }{ 3 }\pi r^2*9=3\pi r^2\] \[\pi r l=3 \pi r^2\] \[\frac{ \pi r l }{ \pi r^2 }=3,\frac{ l }{ r }=3\]

OpenStudy (midhun.madhu1987):

now find the ration of l and r